Consider the accompanyi g data en fuxural strength (MPaforcencrete bams fcertain ty 07 8.1867.0 7 78 6.3 7 7.4 11.6 9.7 63 S3 7.4 72 6.3 78 7,7 70 7 6.8 113 7.3 9.0 18 8.7 Stat& hich ?timator ,ou used nterpret thia point eztimate
Answer #1

we are given the following data:

10.7 8.1 8.6 7.0 5.7 7.8 6.3
7.7 7.4 11.6 9.7 6.3 6.8 7.4
7.2 6.5 7.8 7.7 7.0 7.9 6.8
9.7 11.3 7.3 9.0 11.8 8.7

ANSWER TO :

PART (a) Calculate a point estimate of the mean value of strength for the conceptual population of all beams manufactured in this fashion.

total number of observations, n = 27

We use the sample mean, \bar{x} to estimate the population mean \mu

the point estimate of the mean value of strength for the conceptual population of all beams manufactured in this fashion will be given by:

\bar{x} = sum of all the observations / total number of observations

\bar{x} = \Sigmaxi / n

given, \Sigmaxi = 219.8

therefore, \bar{x} = 219.8 / 27

\bar{x} = 8.141

The estimator we used is the sample mean, i.e., \bar{x}

PART (b) Calculate a point estimate of the strength value that separates the weakest 50% of all such beams from the strongest 50%

we use the sample median.

Recall that, a median is the middle value in a sequence which is sorted in ascending order, i.e., from lowest to highest.

hence, the median will be the value which separates the weakest 50% of beams from the strongest 50%.

to calculate the median first we arrange the data in ascending order

1 - 5.7

2 - 6.3

3 - 6.3

4 - 6.5

5 - 6.8

6 - 6.8

7 - 7.0

8 - 7.0

9 - 7.2

10 - 7.3

11 - 7.4

12 - 7.4

13 - 7.7

14 - 7.7

15 - 7.8

16 - 7.8

17 - 7.9

18 - 8.1

19 - 8.6

20 - 8.7

21 - 9.0

22 - 9.7

23 - 9.7

24 - 10.7

25 - 11.3

26 - 11.6

27 -11.8

here the total no of observations = 27 i.e., odd

the median for odd number of observation is

median, \tilde{x} = [(n+1)/2]th term from the sorted sequence of data

hence,

\tilde{x} = [(27+1)/2]th term from the sorted sequence of data

\tilde{x} = 28/2th term from the sorted sequence of data

\tilde{x} = 14th term   from the sorted sequence of data

1 - 5.7

2 - 6.3

3 - 6.3

4 - 6.5

5 - 6.8

6 - 6.8

7 - 7.0

8 - 7.0

9 - 7.2

10 - 7.3

11 - 7.4

12 - 7.4

13 - 7.7

14 - 7.7 - median

15 - 7.8

16 - 7.8

17 - 7.9

18 - 8.1

19 - 8.6

20 - 8.7

21 - 9.0

22 - 9.7

23 - 9.7

24 - 10.7

25 - 11.3

26 - 11.6

27 -11.8

Hence, the median is 7.7 and the estimator we used is the median of the data i.e., \tilde{x}

PART (c) Calculate and interpret a point estimate of the population standard deviation

teh point estimate of the population standard deviation is the sample standard deviation

the sample standard deviation, s is given by:

s = \sqrt{}[ (\Sigmaxi2 -(\Sigmaxi)2/n)/ (n-1)]

given that,

\Sigmaxi2 = 1862.34

\Sigmaxi = 219.8

n = 27

therefore,

s = \sqrt{}[(1862.34 - (219.8)2/27)/(27-1)]

s = \sqrt{}(1862.34 - 1789.34)/26]

s = \sqrt{}73/26

s = \sqrt{}2.80769 = 1.6756 = 1.676 (rounded off)

hence, the estimate for population standard deviation is 1.676 and the estimator we used is sample

standard deviation, s

we know the standard deviation describes the spread or dispersion of the data.

therefore, interpreting the point estimate, s

The estimate describes the spread of the data.

PART (d) Calculate a point estimate of the proportion of all such beams whose flexural strength exceeds 10 MPa

let the "success" be = all the observations greater than 10 MPa

therefore, the successes are:

10.7

11.3

11.6

11.8

total number of successes = 4

therefore the proportion of all such beams = total number of successes / total number of observations

we know,

total number of successes = 4

total number of observations = 27

therefore, required proportion, \hat{p} = 4/27

\hat{p} = 0.1481 = 0.148 (rounded off)

point estimate of the proportion of all such beams whose flexural strength exceeds 10 MPa is 0.148

PART (e) Calculate a point estimate of the population coefficient of variation \sigma/\mu

the point estimate of population coefficient of variation will be

sample standard deviation / sample mean = s / \bar{x}

s / \bar{x} = 1.676 / 8.141

s / \bar{x} = 0.20587

s / \bar{x} = 0.2059 (rounded off)